. A flywheel used for storing energy consists of a uniform disk of mass 1.5x105
kg and radius 2.2 m that
rotates at 3000 rev/min about its center of mass Find its kinetic energy.
1
Expert's answer
2020-12-06T17:17:59-0500

"Kinetic\space Energy=\frac{1}{2}\times I\times w\times\ w
\\whereby: \space \\
\qquad I=moment \space of \space \space Inertia \\
\qquad w=rotational\space velocity(rpm) 
\\Therefore;\\\qquad I=k\times m\times r \times r \\
 \quad =\frac{1}{2}\times 1.5\times 10^{5} \times\ 2.2\times 2.2\\
\quad = 75000\times 2.2\times 2.2\\
\quad=363000
\\
Thus:\\
Kinetic\space Energy= \frac{1}{2}\times 363000\times\ 3000\times 3000\\
\quad = 1,629,000,000,000Joules
\\ \quad=1.629\times10^{12}Joules."